OpenStudy (anonymous):

Which of the following represents a difference of squares binomial? Answer (k2 – 9)2 k4 + 81 (k – 3)2 k4 – 81

OpenStudy (anonymous):

k^4 - 81 = (k^2 - 9)(k^2 + 9) because k^4 = (k^2) squared